Whats the present value, when interest rates are 75 percent, of a $155 payment made every year forever? (Round decimal places.) your answer to 2
media%2F48b%2F48b55878-3aea-478c-9aea-03
Answer #1

Question 1

This question talks about $155 FOREVER. This implies it is an perpetuity.

Present Value of perpetuity = Annual Payment/Interest Rate

Present Value of Perpetuity = $155/7.5% = $2,066.67 (answer)

Question 2

Effective Annual rate = (1 + Periodic Rate)n - 1

where n is the number of periods in 1 year.

In this question, 1 compounding period is for 2 weeks

Number of Weeks in 1 year = 52

Therefore, number of compunding periods = 52/2 = 26 periods = n

EAR = (1 + 30%)26 - 1

EAR = 917.3333 - 1

EAR = 916.3333

EAR = 91633.33%